Let p(X) be a quadratic polynomial p(3)=9&p(-3)=-10.then the coefficient of X in p(X)?

Grade:10

1 Answers

Arun
6 years ago
Dear student
 
Let f(x) = ax2 + bx + c
Given, f(3) = 9
This gives 9a + 3b + c = 9        ... (1)
 
Given, f(-3) = -10
This gives 9a - 3b + c = -10        ... (2)
 
Subtracting (2) from (1), we get,
6b = 19
b = 19/6
 
Thus, the coefficient of x in f(x) is b = 19/6.
 
Regards
Arun (askIITians forum expert)